OpenStudy (anonymous):

derivate 1-cosx/sinx

OpenStudy (anonymous):

paskaa :D

OpenStudy (alekos):

cosec^2(x)

OpenStudy (mathmath333):

d/dx(1-(cosx/sinx)) =d/dx(1-cotx) =d/dx(1)-d/dx(cotx) =0-(-cosec^2x) =cosec^2x